# LLMEasyQuant: Scalable Quantization for Parallel and Distributed LLM Inference

### v5

@@ A.1.1 Scale Factor Convergence Analysis
where $\hat{X}$ and $\hat{W}$ are the quantized versions of $X$ and $W$ respectively.
Proof of Theorem A.1.
We begin by analyzing the quantization error decomposition. The quantization error can be decomposed as:
$\|XW-\hat{X}\hat{W}\|_{F}^{2}$ $=\|XW-(X/s_{j})(W\cdot s_{j})\|_{F}^{2}$ (14)
$=\|XW-XW\|_{F}^{2}=0$ (15)
@@
For the quantized versions, we analyze the error propagation:
$\|\hat{X}\hat{W}-XW\|_{F}^{2}$ $=\|\text{Quantize}(X/s_{j})\cdot\text{Quantize}(W\cdot s_{j})-XW\|_{F}^{2}$ (17)
@@
Let $\delta_{X}$ and $\delta_{W}$ be the quantization step sizes for activations and weights respectively. Since quantization error is bounded by $\frac{\delta}{2}$ where $\delta$ is the quantization step size, and $\delta\propto\frac{1}{2^{b-1}-1}$, we have:
$\|\text{Quantize}(X/s_{j})-X/s_{j}\|_{F}^{2}\leq\frac{B\cdot D\cdot\delta_{X}^{2}}{4}$ (20)
$\|\text{Quantize}(W\cdot s_{j})-W\cdot s_{j}\|_{F}^{2}\leq\frac{D\cdot D^{\prime}\cdot\delta_{W}^{2}}{4}$ (21)
Therefore:
$\|\hat{X}\hat{W}-XW\|_{F}^{2}\leq\frac{B\cdot D\cdot\delta_{X}^{2}\cdot\|W\cdot s_{j}\|_{F}^{2}}{4}+\frac{D\cdot D^{\prime}\cdot\delta_{W}^{2}\cdot\|X/s_{j}\|_{F}^{2}}{4}$ (22)
As the bitwidth $b$ increases, $\delta_{X},\delta_{W}\to 0$, and thus the quantization error approaches zero, proving the convergence. ∎
A.1.2 Optimal Scale Factor Derivation
We derive the optimal scale factor that minimizes quantization error.
Lemma 1 (Optimal Scale Factor).
The optimal scale factor $s_{j}^{*}$ that minimizes the quantization error is given by:
@@ A.1.2 Optimal Scale Factor Derivation
Proof of Lemma A.1.
We minimize the expected quantization error:
$\mathbb{E}[\|XW-\hat{X}\hat{W}\|_{F}^{2}]$ $=\mathbb{E}\left[\left\|\frac{X}{s_{j}}\cdot(W\cdot s_{j})-\text{Quantize}\left(\frac{X}{s_{j}}\right)\cdot\text{Quantize}(W\cdot s_{j})\right\|_{F}^{2}\right]$ (24)
Using the quantization error bounds and taking the derivative with respect to $s_{j}$:
$\frac{\partial}{\partial s_{j}}\mathbb{E}[\|XW-\hat{X}\hat{W}\|_{F}^{2}]=0$ (25)
This yields:
$s_{j}^{*}=\sqrt{\frac{\mathbb{E}[\max(|X_{j}|)^{2}]}{\mathbb{E}[\max(|W_{j}|)^{2}]}}$ (26)
The SmoothQuant formulation $s_{j}=\left(\frac{\max(|X_{j}|)^{\alpha}}{\max(|W_{j}|)^{1-\alpha}}+\epsilon\right)$ approximates this optimal scale factor with $\alpha=0.5$ providing the best approximation. ∎
A.1.3 Error Bound Analysis for SimQuant
We derive tight error bounds for SimQuant’s dynamic quantization approach through rigorous mathematical analysis.
Theorem 2 (SimQuant Reconstruction Error Bound).
For SimQuant with bitwidth $b$ and channel-wise quantization, the reconstruction error is bounded by:
$\|X-\hat{X}\|_{\infty}\leq\frac{\max(X)-\min(X)}{2^{b}-1}$ (27)
where $\hat{X}$ is the quantized version of $X$.
Proof of Theorem A.2.
@@ A.1.3 Error Bound Analysis for SimQuant
$\|X-\hat{X}\|_{\infty}=\max_{i,d}|X_{i,d}-\hat{X}_{i,d}|\leq\frac{\max(X)-\min(X)}{2^{b}-1}$ (32)
This completes the proof. ∎
A.1.4 Convergence Rate Analysis for SimQuant
We establish the convergence rate of SimQuant’s dynamic quantization.
Lemma 2 (SimQuant Convergence Rate).
@@ A.1.4 Convergence Rate Analysis for SimQuant
Proof of Lemma A.2.
The convergence rate analysis follows from the error bound in Theorem A.2. As the bitwidth $b$ increases, the quantization step size decreases exponentially:
$\Delta_{d}=\frac{v_{\max}^{(d)}-v_{\min}^{(d)}}{2^{b}-1}=O\left(\frac{1}{2^{b}}\right)$ (33)
Therefore, the reconstruction error decreases as:
$\|X-\hat{X}\|_{\infty}\leq\frac{\max(X)-\min(X)}{2^{b}-1}=O\left(\frac{1}{2^{b}}\right)$ (34)
This establishes the exponential convergence rate with respect to bitwidth. ∎
A.2 Optimization Guarantees for Layer-wise Quantization
A.2.1 Mixed-Precision Search Convergence Analysis
We establish rigorous convergence guarantees for the mixed-precision bitwidth search algorithm through detailed mathematical analysis.
Theorem 3 (Mixed-Precision Search Convergence).
@@ A.2.1 Mixed-Precision Search Convergence Analysis
Proof of Theorem A.3.
We analyze the convergence properties of the mixed-precision search algorithm step by step.
Step 1: Finite Search Space Analysis The search space $\mathcal{B}=\{2,3,4,8\}$ is finite with cardinality $|\mathcal{B}|=4$. For a model with $L$ layers, the total search space has cardinality $|\mathcal{B}|^{L}$, which is finite.
Step 2: Objective Function Properties The objective function $f(\{b_{\ell}\})=\mathcal{L}_{\text{task}}+\lambda\sum_{\ell}\Phi(b_{\ell})$ has the following properties:
  • $\mathcal{L}_{\text{task}}\geq 0$ (loss functions are non-negative)
  • $\Phi(b_{\ell})\geq 0$ (cost functions are non-negative)
  • $f(\{b_{\ell}\})\geq 0$ (bounded below by zero)
Step 3: Greedy Optimization Convergence The algorithm performs greedy optimization by iteratively updating each layer’s bitwidth:
$b_{\ell}^{(t+1)}=\arg\min_{b\in\mathcal{B}}f(b_{1}^{(t)},\ldots,b_{\ell-1}^{(t)},b,b_{\ell+1}^{(t)},\ldots,b_{L}^{(t)})$ (36)
Since the objective function is bounded below and the search space is finite, the sequence $\{f(\{b_{\ell}^{(t)}\})\}_{t=0}^{\infty}$ is monotonically decreasing and bounded below. By the monotone convergence theorem, this sequence converges to a limit.
Step 4: Convergence to Local Optimum The algorithm terminates when no single-layer bitwidth change can improve the objective function. This corresponds to reaching a local optimum where:
$f(\{b_{\ell}^{*}\})\leq f(\{b_{\ell}\})\quad\forall\{b_{\ell}\}\text{ in the neighborhood of }\{b_{\ell}^{*}\}$ (37)
Step 5: Convergence Rate Analysis The convergence rate is $O(L\cdot|\mathcal{B}|)$ because:
  • Each iteration evaluates at most $L\cdot|\mathcal{B}|$ configurations
  • The algorithm terminates when no improvement is found
  • The worst-case number of iterations is bounded by the total number of possible configurations
This completes the proof. ∎
A.2.2 Distributed Quantization Synchronization Analysis
We prove the correctness and efficiency of distributed quantization parameter synchronization.
Theorem 4 (Distributed Synchronization Correctness).
@@ A.2.2 Distributed Quantization Synchronization Analysis
Proof of Theorem A.4.
We analyze the distributed synchronization mechanism step by step.
Step 1: NCCL AllGather Properties The NCCL AllGather operation has the following properties:
  • Deterministic: Given identical inputs, AllGather produces identical outputs
  • Collective: All processes participate in the operation
  • Atomic: The operation completes on all processes simultaneously
Step 2: Parameter Synchronization For quantization parameters $\delta_{\ell}^{(p)}$ and $z_{\ell}^{(p)}$ computed on device $p$:

### v6

@@ B.2.1 Preliminary Lemmas
where $Q(\cdot)$ denotes the quantization operator and $\hat{X}$, $\hat{W}$ are the quantized versions of $X/s_{j}$ and $W\cdot s_{j}$ respectively.
Proof B.8 (Proof of Lemma A.1).
The algebraic equivalence $(X/s_{j})\cdot(W\cdot s_{j})=X\cdot W$ ensures that before quantization, the transformation preserves the original matrix multiplication. The quantization error arises solely from the quantization operators $Q(\cdot)$ applied to the scaled tensors, leading to the stated decomposition.
Lemma B.9 (Bound on Quantization Operator).
There exists an absolute constant $c>0$ such that, for any tensor $Z\in\mathbb{R}^{m\times n}$ with quantization step size $\delta=\frac{2\max(|Z|)}{2^{b}-1}$, the quantization error satisfies:
$\|Q(Z)-Z\|_{F}^{2}\leq c\cdot\frac{mn\cdot\max(|Z|)^{2}}{(2^{b}-1)^{2}}$ (22)
where $b$ is the quantization bitwidth.
Proof B.10 (Proof of Lemma A.2).
For each element $Z_{i,j}$, the quantization error is bounded by half the quantization step size:
$|Q(Z_{i,j})-Z_{i,j}|\leq\frac{\delta}{2}=\frac{\max(|Z|)}{2^{b}-1}$ (23)
Taking the Frobenius norm over all $mn$ elements gives the stated bound.
B.2.2 Scale Factor Convergence Analysis
Theorem B.11 (SmoothQuant Scale Factor Convergence).
There exists an absolute constant $c>0$ such that, for any $\epsilon\in(0,1)$, if we choose the SmoothQuant scale factor $s_{j}=\left(\frac{\max(|X_{j}|)^{\alpha}}{\max(|W_{j}|)^{1-\alpha}}+\epsilon\right)$ with $\alpha\in[0,1]$, then for activation tensors $X\in\mathbb{R}^{B\times D}$ and weight matrices $W\in\mathbb{R}^{D\times D^{\prime}}$, the quantization error satisfies:
$\mathbb{E}[\|XW-\hat{X}\hat{W}\|_{F}^{2}]\leq c\cdot\frac{\max(|X_{j}|)^{2}+\max(|W_{j}|)^{2}\cdot s_{j}^{2}}{s_{j}^{2}\cdot(2^{b}-1)^{2}}\cdot BD\cdot DD^{\prime}$ (24)
where $\hat{X}$ and $\hat{W}$ are the quantized versions of $X$ and $W$ respectively, and $b$ is the quantization bitwidth. In particular, as $b\to\infty$, we have $\lim_{b\to\infty}\mathbb{E}[\|XW-\hat{X}\hat{W}\|_{F}^{2}]=0$.
Proof B.12 (Proof of Theorem A.1).
We prove this theorem step by step, using the lemmas established above.
Step 1: Error Decomposition
By Lemma B.7, we have:
$\|XW-\hat{X}\hat{W}\|_{F}^{2}=\|Q(X/s_{j})Q(W\cdot s_{j})-(X/s_{j})(W\cdot s_{j})\|_{F}^{2}$ (25)
@@ B.2.2 Scale Factor Convergence Analysis
Step 2: Triangle Inequality Application
For the quantized versions, we analyze the error propagation using the triangle inequality:
$\|\hat{X}\hat{W}-XW\|_{F}^{2}$ $=\|\text{Quantize}(X/s_{j})\cdot\text{Quantize}(W\cdot s_{j})-(X/s_{j})(W\cdot s_{j})\|_{F}^{2}$ (27)
@@
Step 3: Quantization Error Bounds
Let $\delta_{X}$ and $\delta_{W}$ be the quantization step sizes for activations and weights respectively. By Lemma B.9, we have:
$\|\text{Quantize}(X/s_{j})-X/s_{j}\|_{F}^{2}\leq c\cdot\frac{B\cdot D\cdot\max(|X/s_{j}|)^{2}}{(2^{b}-1)^{2}}$ (30)
$\|\text{Quantize}(W\cdot s_{j})-W\cdot s_{j}\|_{F}^{2}\leq c\cdot\frac{D\cdot D^{\prime}\cdot\max(|W\cdot s_{j}|)^{2}}{(2^{b}-1)^{2}}$ (31)
where $\delta_{X}=\frac{2\max(|X/s_{j}|)}{2^{b}-1}$ and $\delta_{W}=\frac{2\max(|W\cdot s_{j}|)}{2^{b}-1}$.
Step 4: Final Bound
Combining the bounds, we obtain:
$\|\hat{X}\hat{W}-XW\|_{F}^{2}\leq c\cdot\frac{\max(|X/s_{j}|)^{2}\cdot\|W\cdot s_{j}\|_{F}^{2}+\max(|W\cdot s_{j}|)^{2}\cdot\|X/s_{j}\|_{F}^{2}}{(2^{b}-1)^{2}}\cdot BD\cdot DD^{\prime}$ (32)
As the bitwidth $b$ increases, $\delta_{X},\delta_{W}\to 0$, and thus the quantization error approaches zero, proving the convergence. This completes the proof.
B.2.3 Optimal Scale Factor Derivation
Lemma B.13 (Optimal Scale Factor).
The optimal scale factor minimizing quantization error is:
@@ B.2.3 Optimal Scale Factor Derivation
Proof B.14 (Proof of Lemma A.1).
Minimize: $\mathcal{L}(s_{j})=\mathbb{E}[\|XW-Q(X/s_{j})Q(W\cdot s_{j})\|_{F}^{2}]$.
Using error bounds: $\mathcal{L}(s_{j})\approx\mathbb{E}\left[\frac{BD\delta_{X}^{2}\|W\cdot s_{j}\|_{F}^{2}}{4}+\frac{DD^{\prime}\delta_{W}^{2}\|X/s_{j}\|_{F}^{2}}{4}\right]$.
Substituting $\delta_{X}=\frac{2\max(|X/s_{j}|)}{2^{b}-1}$, $\delta_{W}=\frac{2\max(|W\cdot s_{j}|)}{2^{b}-1}$:
$\mathcal{L}(s_{j})$ $\propto\mathbb{E}\left[\frac{\max(|X/s_{j}|)^{2}\|W\cdot s_{j}\|_{F}^{2}}{s_{j}^{2}}+\frac{\max(|W\cdot s_{j}|)^{2}\|X/s_{j}\|_{F}^{2}}{s_{j}^{2}}\right]$ (34)
$\propto\mathbb{E}\left[\frac{\max(|X_{j}|)^{2}}{s_{j}^{2}}+s_{j}^{2}\max(|W_{j}|)^{2}\right]$ (35)
Taking derivative: $\frac{\partial\mathcal{L}}{\partial s_{j}}=-\frac{2\mathbb{E}[\max(|X_{j}|)^{2}]}{s_{j}^{3}}+2s_{j}\mathbb{E}[\max(|W_{j}|)^{2}]=0$.
Solving: $s_{j}^{*}=\sqrt{\frac{\mathbb{E}[\max(|X_{j}|)^{2}]}{\mathbb{E}[\max(|W_{j}|)^{2}]}}$.
SmoothQuant approximation: $s_{j}=\left(\frac{\max(|X_{j}|)^{\alpha}}{\max(|W_{j}|)^{1-\alpha}}\right)$ with $\alpha=0.5$ minimizes approximation error.
B.2.4 Error Bound Analysis for SimQuant
Theorem B.15 (SimQuant Reconstruction Error Bound).
There exists an absolute constant $c>0$ such that, for any $\epsilon\in(0,1)$, if we apply SimQuant with bitwidth $b$ and channel-wise quantization to tensor $X\in\mathbb{R}^{B\times D}$, then with probability at least $1-\epsilon$, the reconstruction error is bounded by:
$\|X-\hat{X}\|_{\infty}\leq c\cdot\frac{\max_{d\in[D]}(\max_{i}X_{i,d}-\min_{i}X_{i,d})}{2^{b}-1}$ (36)
where $\hat{X}$ is the quantized version of $X$, $B$ is the batch size, and $D$ is the feature dimension.
Proof B.16 (Proof of Theorem A.2).
@@ B.2.4 Error Bound Analysis for SimQuant
$\|X-\hat{X}\|_{\infty}=\max_{i,d}|X_{i,d}-\hat{X}_{i,d}|\leq c\cdot\frac{\max_{d\in[D]}(\max_{i}X_{i,d}-\min_{i}X_{i,d})}{2^{b}-1}$ (41)
where $c$ is an absolute constant. This completes the proof.
B.2.5 Convergence Rate Analysis for SimQuant
Lemma B.17 (SimQuant Convergence Rate).
@@ B.2.5 Convergence Rate Analysis for SimQuant
Proof B.18 (Proof of Lemma A.2).
From Theorem A.2: $\|X-\hat{X}\|_{\infty}\leq\frac{\max(X)-\min(X)}{2^{b}-1}$.
As $b\to\infty$: $\Delta_{d}=\frac{v_{\max}^{(d)}-v_{\min}^{(d)}}{2^{b}-1}=O(2^{-b})$.
Therefore: $\|X-\hat{X}\|_{\infty}=O(2^{-b})$, establishing exponential convergence.
B.3 Optimization Guarantees for Layer-wise Quantization
B.3.1 Mixed-Precision Search Convergence Analysis
Theorem B.19 (Mixed-Precision Search Convergence).
@@ B.3.1 Mixed-Precision Search Convergence Analysis
Proof B.20 (Proof of Theorem A.3).
Search space: $\mathcal{B}=\{2,3,4,8\}$, $|\mathcal{B}|=4$, total space: $|\mathcal{B}|^{L}$.
Objective: $f(\{b_{\ell}\})=\mathcal{L}_{\text{task}}+\lambda\sum_{\ell=1}^{L}\Phi(b_{\ell})$ where $\mathcal{L}_{\text{task}}\geq 0$, $\Phi(b_{\ell})\geq 0$, hence $f\geq 0$.
Greedy update: $b_{\ell}^{(t+1)}=\arg\min_{b\in\mathcal{B}}f(b_{1}^{(t)},\ldots,b_{\ell-1}^{(t)},b,b_{\ell+1}^{(t)},\ldots,b_{L}^{(t)})$.
Sequence $\{f^{(t)}\}$ is monotone decreasing: $f^{(t+1)}\leq f^{(t)}$ and bounded: $f^{(t)}\geq 0$.
By monotone convergence: $\lim_{t\to\infty}f^{(t)}=f^{*}$ exists.
Termination condition: $\forall\ell,\forall b\in\mathcal{B}$:
$f(b_{1}^{*},\ldots,b_{\ell-1}^{*},b_{\ell}^{*},b_{\ell+1}^{*},\ldots,b_{L}^{*})\leq f(b_{1}^{*},\ldots,b_{\ell-1}^{*},b,b_{\ell+1}^{*},\ldots,b_{L}^{*})$ (43)
This defines local optimum: $f(\{b_{\ell}^{*}\})\leq f(\{b_{\ell}\})$ for all $\{b_{\ell}\}$ in neighborhood.
Complexity: each iteration evaluates $\leq L\cdot|\mathcal{B}|$ configurations, worst-case iterations $\leq|\mathcal{B}|^{L}$, hence $T=O(L\cdot|\mathcal{B}|\cdot|\mathcal{B}|^{L})=O(L\cdot|\mathcal{B}|^{L+1})$.
B.3.2 Distributed Quantization Synchronization Analysis
Theorem B.21 (Distributed Synchronization Correctness).
@@ B.3.2 Distributed Quantization Synchronization Analysis
Proof B.22 (Proof of Theorem A.4).
AllGather properties: deterministic ($\text{AllGather}(x_{1},\ldots,x_{P})=\text{AllGather}(x_{1}^{\prime},\ldots,x_{P}^{\prime})$ if $x_{p}=x_{p}^{\prime}$), collective (all $P$ processes participate), atomic (completes simultaneously).
Synchronization: $\delta_{\ell}^{\text{global}}=\text{AllGather}(\delta_{\ell}^{(1)},\ldots,\delta_{\ell}^{(P)})$, $z_{\ell}^{\text{global}}=\text{AllGather}(z_{\ell}^{(1)},\ldots,z_{\ell}^{(P)})$.
By determinism: $\delta_{\ell}^{(p)}=\delta_{\ell}^{\text{global}}$, $z_{\ell}^{(p)}=z_{\ell}^{\text{global}}$ for all $p\in[P]$.
Quantized weights: $\hat{W}_{\ell}^{(p)}=Q(W_{\ell}^{(p)},\delta_{\ell}^{\text{global}},z_{\ell}^{\text{global}})$.
Since $Q$ is deterministic: $\hat{W}_{\ell}^{(p)}=\hat{W}_{\ell}^{\text{global}}$ for all $p$, ensuring consistency.
B.4 Computational Complexity Analysis
B.4.1 Algorithmic Complexity
Theorem B.23 (SmoothQuant Complexity).
Proof B.24 (Proof of Complexity).
Operations:
$T_{\text{scale}}$ $=O(D)+O(DD^{\prime})=O(DD^{\prime})$ (44)
$T_{\text{smooth}}$ $=O(BD)$ (45)
$T_{\text{quant}}$ $=O(BD+DD^{\prime})$ (46)
Total time: $T=O(BD+DD^{\prime})=O(BDD^{\prime})$ (dominated by GEMM).
Space: $S=O(BD+DD^{\prime})$.
B.4.2 Memory Hierarchy Optimization
Theorem B.25 (Memory Bandwidth Optimization).
Proof B.26 (Proof of Memory Optimization).
For separate operations, the memory bandwidth requirements include loading FP16 weights ($2\times|W|$ bytes), storing quantized weights ($b/8\times|W|$ bytes), and loading quantized weights for GEMM ($b/8\times|W|$ bytes), resulting in a total of $(2+2\times b/8)\times|W|$ bytes.
For fused operation, the memory bandwidth requirements include loading FP16 weights ($2\times|W|$ bytes) and storing quantized weights ($b/8\times|W|$ bytes), resulting in a total of $(2+b/8)\times|W|$ bytes.
B.5 Error Propagation Analysis
B.5.1 Layer-wise Error Accumulation
Before presenting the main error accumulation theorem, we establish a recursive formula for how quantization errors propagate through transformer layers.
Lemma B.27 (Recursive Error Propagation).
For a transformer with $L$ layers, let $f_{\ell}$ denote the function at layer $\ell$ and $\hat{f}_{\ell}$ denote its quantized version. If $\epsilon_{\ell}$ is the quantization error at layer $\ell$, then the accumulated error through the network satisfies:
where $J_{j}=\frac{\partial f_{j}}{\partial x}$ is the Jacobian of layer $j$ at the input point.
Proof B.28 (Proof of Lemma A.3).
We prove this by induction on the number of layers. For $L=1$, the statement is trivial. For $L>1$, we use the chain rule and the fact that quantization errors are bounded:
$\|f_{L}\circ\cdots\circ f_{1}(x)-\hat{f}_{L}\circ\cdots\circ\hat{f}_{1}(x)\|$ $\leq\|f_{L}\circ\cdots\circ f_{1}(x)-f_{L}\circ\cdots\circ f_{2}\circ\hat{f}_{1}(x)\|$ (48)
$\quad+\|f_{L}\circ\cdots\circ f_{2}\circ\hat{f}_{1}(x)-\hat{f}_{L}\circ\cdots\circ\hat{f}_{1}(x)\|$ (49)
The first term is bounded by $\epsilon_{1}\prod_{j=2}^{L}\|J_{j}\|$, and the second term follows by the inductive hypothesis.
Theorem B.29 (Error Accumulation Bound).
Proof B.30 (Proof of Error Accumulation).
By Lemma B.27, we have:
where $J_{j}=\frac{\partial f_{j}}{\partial x}$ is the Jacobian of layer $j$ and $\hat{f}_{\ell}$ is the quantized version of layer $\ell$.
Since $\epsilon_{\ell}\leq\epsilon$ for all layers, the accumulated error is $O(L\cdot\epsilon)$. This completes the proof.
B.6 Calibration Data Requirements Analysis
B.6.1 Minimum Calibration Set Size
Theorem B.31 (Calibration Data Requirements).
Proof B.32 (Proof of Calibration Requirements).
Hoeffding’s bound for scale estimation: $P(|\hat{s}-s|\geq\epsilon)\leq 2\exp\left(-\frac{2n\epsilon^{2}}{(b-a)^{2}}\right)$ where $n$ is sample size, $[a,b]$ is data range.
For $D$ dimensions, union bound:
$P(\exists j:|\hat{s}_{j}-s_{j}|\geq\epsilon)\leq\sum_{j=1}^{D}P(|\hat{s}_{j}-s_{j}|\geq\epsilon)\leq D\cdot 2\exp\left(-\frac{2n\epsilon^{2}}{(b-a)^{2}}\right)$ (52)
Setting $D\cdot 2\exp\left(-\frac{2n\epsilon^{2}}{(b-a)^{2}}\right)=\delta$ and solving:
$n$ $\geq\frac{(b-a)^{2}}{2\epsilon^{2}}\log\left(\frac{2D}{\delta}\right)$ (53)
$=O\left(\frac{D\log D}{\epsilon^{2}}\right)$ (54)
Lemma B.33.
Proof B.34.